1. Valles de Benasque y Eriste

MIRADOR DEL ’PICHIVILLO’ (PICHIVILLO BELLEVUE, CERLER) 1 hour

Take the path opposite the ”Ribagorza building” and ”La Parrilla bar” on the other side of the road on the direction to the graveyard. You will pass next to it after a 20 min. walk. Then the path crosses the Regoso ravine and goes into the wood. When you get to a deviation, turn left and after about 6 minutes of steep ascension you will get to the Mirador del Pichivillo (Pichivillo Bellevue), a high rock with beautiful, stunning views.